A cylindrical tube open at both ends has been made of iron sheet which is 2cm thick. If the outer diameter is 16 cm and its length is 100 cm, find how many cubic centimeters of iron has been used in making the tube.

A  cylindrical  tube  open at  both ends has been made of iron sheet  which is 2cm  thick.  If the outer  diameter is 16 cm and its length is 100 cm, find  how many cubic centimeters of iron has been  used in making the tube.

Grade:9

2 Answers

Ravi
9 years ago
Outer dia=16
Outer radius= 8cm= Router

Inner radius= Outer radius – thickness of tube= 8-2= 6 cm

So, Volume of iron used= pi*h*(R2outer-R2inner)= Volume of outer cylinder- volume of inner cylinder
Nishant Vora IIT Patna
9 years ago
Do=16cm
thickness t = 2 cm
So Inner dia Di=16-4 = 12cm
Length = 100cm

So outer radius Ro=8cm
So inner radius Ri=6cm

So volume of tube = \piRo2h –\pi Ri2h
Now put all the values